Write the linear equation satisfying each of the following conditions: parallel to the line 2x - 3y = 6 and has a y-intercept of 7. [Write the equation in the form ax + by + c = 0].
Answer by palanisamy(496)

The given equation of the line is 2x-3y = 6
The equation of a line parallel to this line is
2x-3y+c = 0 ....(1) where c is a constant.
Given, the y-intercept is 7.
That is the point (0,7) lies on the line (1).
Put x\0 and y=7 in (1)
We get, 2*0-3*7+c = 0
-21+c=0
c=21
Therefore (1)=>, the equation of the required line is 2x-3y+21 = 0